OpenStudy (anonymous):

Calculate the volume of a sphere where the radius is 9 meters A. 81 cubic meters B. 324 cubic meters C. 972 cubic meters D. 2916 cubic meters

OpenStudy (anonymous):

Volume = 4/3.pi.r^3. Just plug it it and use calculator.

OpenStudy (anonymous):

cant use a calculator

OpenStudy (anonymous):

V = 4/3.pi.9^3 = 4/3.pi.729 = 4/3.pi.243 = 972.pi

OpenStudy (amistre64):

paper and pencil are just as effective :)